OpenStudy (anonymous):

Part 1: Find the perimeter of a rectangular object which has radical 180 a length of radical 125 feet and a width of feet. (3 Points) Part 2: Explain, in complete sentences, how you arrived at the answer and give the final solution in simplified radical form. (2 points) Part 3: What type of object in your home or school might this be? (1 point)

OpenStudy (turingtest):

P=2(l+w) l=180 w=125 plug in then the numbers

OpenStudy (anonymous):

Part 1: √180 = √(36*5) = √36 * √5 = 6√5 √125 = √(25*5) = √25 * √5 = 5√5 perimeter of a rectangular object = 2*√180 + 2*√125 = 2*6√5 + 2*5√5 = 12√5 + 10√5 = 22√5.

OpenStudy (anonymous):

Part 2: final solution in simplified radical form = 22√5 Part 3: √180 is about 13.4 feet √125 is about 11.2 feet